# Convergent Primal-Dual Plug-and-Play Image Restoration: A General Algorithm and Applications

### v2

@@ III-B1 Image Restoration under Gaussian Noise
[…]
@@
1:$\mathbf{u}_{0},\,\mathbf{w}_{0}^{(1)},\,\mathbf{w}_{0}^{(2)},\,\gamma_{1}>0,\,\gamma_{2}>0,\,\varepsilon>0$
@@
4:     $\mathbf{w}_{n}^{(1)}\leftarrow\mathbf{w}_{n}^{(1)}+\gamma_{2}\mathbf{\Phi}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
5:     $\mathbf{w}_{n}^{(2)}\leftarrow\mathbf{w}_{n}^{(2)}+\gamma_{2}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
6:     $\mathbf{w}_{n+1}^{(1)}\leftarrow\mathbf{w}_{n}^{(1)}-\gamma_{2}P_{B^{\mathbf{v}}_{2,\,\varepsilon}}(\frac{1}{\gamma_{2}}\mathbf{w}_{n}^{(1)})$;
@@ III-B2 Image Restoration under Poisson Noise
In this observation model, the data-fedelity is expressed as the GKL divergence (refer to [64]), and image restoration can be reduced to the following optimization problem:
[…]
@@
1:$\mathbf{u}_{0},\,\mathbf{w}_{0}^{(1)},\,\mathbf{w}_{0}^{(2)},\,\gamma_{1}>0,\,\gamma_{2}>0,\,\lambda>0$
@@
4:     $\mathbf{w}_{n}^{(1)}\leftarrow\mathbf{w}_{n}^{(1)}+\gamma_{2}\mathbf{\Phi}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
5:     $\mathbf{w}_{n}^{(2)}\leftarrow\mathbf{w}_{n}^{(2)}+\gamma_{2}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
6:     $\mathbf{w}_{n+1}^{(1)}\leftarrow\mathbf{w}_{n}^{(1)}-\gamma_{2}\mathrm{prox}_{\frac{\lambda}{{\gamma_{2}}}\mathrm{GKL}}(\frac{1}{\gamma_{2}}\mathbf{w}_{n}^{(1)})$;
@@ III-C1 Convergent PnP-FBS [17]
$\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\frac{\lambda}{2}\|\mathbf{\Phi\hat{u}}-\mathbf{v}\|_{2}^{2}+\mathcal{R}_{J}(\mathbf{\hat{u}}),$ (38)
@@ III-C2 Convergent PnP-ADMM [24, 23]
Several works have proposed PnP algorithms based on ADMM (PnP-ADMM) and established their convergence guarantees. Nair et al. have proposed a convergent PnP-ADMM and provided extensive analysis on its fixed-point [23]. Sun et al. have proposed a convergent PnP-ADMM algorithm that handles multiple data-fidelity terms, making it practical for large scale datasets [24]. In both studies, the convergence properties are ensured under the firmly nonexpansiveness of the denoiser. PnP-ADMM can address a wider range of tasks than PnP-FBS, including image restoration with Poisson noise.
@@ III-C3 Convergent PnP-PDS[26]
We show the relationship between this algorithm and our PnP-PDS algorithm in Fig 1. In essence, the algorithm in [26] is a special case of our Algorithm 1 without the box constraint.
@@ III-C4 Regularization by Denoising (RED)[44, 51]
$\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\lambda\|\mathbf{\Phi}\mathbf{\hat{u}}-\mathbf{v}\|^{2}_{2}+\mathbf{\hat{u}}^{\top}(\mathbf{\hat{u}}-\mathfrak{G}(\mathbf{\hat{u}})),$ (43)
$\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\lambda\mathrm{GKL}_{\mathbf{v}}(\hat{\mathbf{u}})+\mathbf{\hat{u}}^{\top}(\mathbf{\hat{u}}-\mathfrak{G}(\mathbf{\hat{u}})),$ (44)
@@
$\nabla(\mathbf{\hat{u}}^{\top}(\mathbf{\hat{u}}-\mathfrak{G}(\mathbf{\hat{u}})))=\hat{\mathbf{u}}-\mathfrak{G}(\mathbf{\hat{u}}).$ (45)
@@
However, RED also has a considerable difficulty: the assumptions for the denoiser. First, local homogeneity is an unrealistic property, especially for deep denoisers. In fact, the gradients of the regularization terms computed in the RED algorithms often contain substantial computational errors due to the lack of local homogeneity in the denoisers [51]. Second, cyclically firmly nonexpansiveness [45] is even stronger assumption than the firmly nonexpansiveness assumed in our work. These challenges make it difficult to ensure the convergence of RED algorithms when we employ state-of-the-art denoisers, including DnCNN.
@@ IV Experiments
[…]
@@ IV-A Experimental Setup
In the two experiments, we considered two types of observation operators, one for blur and the other for random sampling. For the blur case, the operator $\mathbf{\Phi}$ was defined as a circular convolution with a $19\times 19$ kernel size [66, kernel 1]. This kernel was normalized so that $\|\mathbf{\Phi}\|_{\mathrm{op}}$ was equal to $1$. In the case of random sampling, pixels of the image were randomly selected and masked. The random sampling rate was set to $80\%$ for all cases. The operator norm also satisfies $\|\mathbf{\Phi}\|_{\mathrm{op}}=1$.
We evaluated the restoration performance by Peak Signal to Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM). To investigate stability, the update rate $c_{n}$ for $\0n\in\mathbb{N}\backslash\{0\}$ was defined by $c_{n}:={\|\mathbf{u}_{n}-\mathbf{u}_{n-1}\|_{2}}/{\|\mathbf{u}_{n-1}\|_{2}},$ where $\mathbf{u}_{n}$ is the intermediate solution at the $n$-th iteration by each algorithm.
@@
Method Algorithm Regularization
PnP-FBS [17] FBS FNE* DnCNN
PnP-ADMM [24] ADMM FNE* DnCNN
PnP-PDS (Unstable) [25] PDS DnCNN (not FNE*)
PnP-PDS (w/o a box const.) [26] PDS FNE* DnCNN
TV [61] PDS TV
RED (Gaussian) [44] SD FNE* DnCNN
RED (Poisson) [44] PDS FNE* DnCNN
DRUNet [37] HQS DRUNet
BPG [50] BPG Bregman Score Denoiser
Proposed PDS FNE* DnCNN
*FNE: firmly nonexpansive   SD: steepest descent
HQS: half-quadratic splitting
In each experiment, we selected the appropriate methods from PnP-FBS [17], PnP-ADMM [24], PnP-PDS (Unstable) [25], PnP-PDS (w/o a box const.) [26], TV [61], RED [44], DRUNet [37], and Bregman Proximal Gradient (BPG) [50], and compared them with the proposed method. The optimization algorithms and regularization for each method are summarized in Table II. Details of PnP-FBS, PnP-ADMM, and RED are given in Section III-C. For BPG, we followed the stopping criterion in the implementation provided by the authors [50]. For all other methods, the number of iterations was fixed at 1200 for the blur case and 3000 for the random sampling case.
For fair comparison, unless otherwise specified, the same DnCNN distributed by Pesquet et al. [17] was used for regularization in PnP-FBS, PnP-ADMM, PnP-PDS (w/o a box const.), RED, and the proposed method. This denoiser was a firmly nonexpansive DnCNN trained on Gaussian noise with a standard deviation of $0.01$. See [17] for details of the training.
@@ IV-B1 Versatility of a Once-Trained Denoiser
The theoretical convergence guarantees in this work build upon PnP-FBS. Hence, we experimentally confirm that the versatility of the denoiser is enhanced by extending PnP-FBS to PnP-PDS.
[…]
Fig. 2 shows the variation in restoration performance across different $\alpha$ and $\zeta$ values. While both methods exhibit comparable peak performance at $\sigma=0.02$ and $\sigma=0.04$, PnP-FBS diverges at lower noise levels (i.e., $\sigma=0.0025$, $0.005$, and $0.01$), as discussed in Section III-C1. For example, in the case of $\sigma=0.01$, we obtain $\lambda_{\mathrm{opt}}=2.22$ from (40) since $0.00452\sigma\|h\|=$ 0.0045 in this situation. In addition, the convergence condition (41) requires $\lambda<2$ since $\|\mathbf{\Phi}\|_{\mathrm{op}}^{2}=1$, which implies $\zeta<0.9$. In fact, stable performance is not observed for $\zeta\geq 0.9$ when $\sigma=0.01$.
In contrast, the proposed method exhibits reasonable performance across all noise levels. This consistency demonstrates the high versatility of the denoiser trained on the specific noise level (see also the discussion in Section III-C1). We also observed that the sensitivity of performance with respect to $\alpha$ tended to increase with higher noise levels, and that the optimal value of $\alpha$ became closer to $1$ as the noise level increased.
@@ IV-B2 Restoration Performance
Deblurring $0.820$ $0.860$ $0.920$ $0.960$ $1.000$
Inpainting $0.900$ $0.820$ $0.820$ $1.000$ $1.000$
TABLE III: The Values of $\alpha$ Used in Each Setting for the Proposed Method.
@@
PnP-PDS (w/o a box const.) [26] $\underline{37.23}$ $\underline{34.03}$ $\underline{31.00}$ $28.24$ $\mathbf{26.00}$-- - --- - --- - --- - --- -
@@
RED [44] $34.11$ $33.38$ $30.89$ $\mathbf{28.30}$ $24.87$-- - --- - --- - --- - --- -
DRUNet [37] $32.60$ $31.80$ $28.54$ $-8.04$ $-9.41$ $33.21$ $32.68$ $31.36$ $26.71$ $-15.9$
PnP-PDS (Proposed) $\mathbf{37.24}$ $\mathbf{34.04}$ $\mathbf{31.01}$ $\underline{28.26}$ $\mathbf{26.00}$ $\mathbf{35.61}$ $\mathbf{35.25}$ $\mathbf{34.43}$ $\mathbf{33.04}$ $\mathbf{31.00}$
TABLE IV: The Average PSNR [dB] Values for Deblurring and Inpainting on the 7 Images From ImageNet the Best Value is Highlighted in Bold and the Second Best Value is Underlined.
Figure 3: The evolution of $\|\nabla Q(\mathbf{x}_{n})\|_{\mathrm{sp}}^{2}$ of PnP-PDS (w/o a box constraint), PnP-FBS, RED, and the proposed method for inpainting at $\sigma=0.01$. The values in each plot are computed as the mean over 7 images.
Figure 4: The evolution of PSNR [dB] and $c_{n}$ of PnP-PDS (Unstable) and the proposed method over iterations for deblurring at $\sigma=0.01$. Each plot in the graph corresponds to each image extracted from ImageNet [65].
@@
We compare the performance of the proposed method with other methods, namely PnP-FBS[17], PnP-PDS (Unstable)[25], PnP-PDS (w/o a box const.) [26], TV[61], RED[44], and DRUNet [37]. For constrained formulations such as PnP-PDS (Unstable), PnP-PDS (w/o a box const.), TV, and the proposed method, we set $\varepsilon=\alpha\sigma\sqrt{K}$ in the same way as in the previous section. For each of these methods, we performed a linear search for $\alpha$ to evaluate the restoration performance. In particular, we provide the values of $\alpha$ for the proposed method in Table III. For existing methods, we adopted the recommended values if parameter settings were suggested by the authors. Otherwise, we performed a linear search over the balancing parameter.
@@
Table IV shows the average PSNR values over seven images, obtained with the optimal parameters for each method. The proposed method demonstrates superior performance compared to PnP-FBS, even though both methods employ the same denoiser for regularization. This is attributed to the difference in the versatility of the denoiser, as mentioned in Section IV-B1.
Furthermore, in the case of inpainting by PnP-FBS, PnP-PDS (w/o a box const.), and RED, the algorithms do not converge and PSNR values are not obtained. To discuss this point, let us recall that the denoiser was trained with the penalty in (17), which was designed to promote nonexpansiveness at low-noise images sampled according to (18). During the PnP iterations, however, the inputs to the denoiser can be much more noisy than those images. In such cases, the denoiser may not maintain its firmly nonexpansiveness, which can further amplify instability through successive iterations. For a numerical verification, we present the evolution of $\|\nabla Q(\mathbf{x}_{n})\|_{\mathrm{sp}}^{2}$ at each iteration in Fig. 3. Although PnP-PDS (without a box constraint), PnP-FBS, RED, and the proposed method employ the same denoiser, the values of $\|\nabla Q(\mathbf{x}_{n})\|_{\mathrm{sp}}^{2}$ in PnP-PDS (without a box constraint), PnP-FBS, and RED exhibit unstable behavior and tend to diverge. In contrast, the proposed method consistently maintains values of $\|\nabla Q(\mathbf{x}_{n})\|_{\mathrm{sp}}^{2}$ that satisfy the condition given in (16). While these observations are numerical, they suggest that the denoiser behaves as a firmly nonexpansive operator during the PnP iterations, thanks to the box constraint.

### v5

@@ III-B1 Image Restoration under Gaussian Noise
$\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\mathcal{R}_{J}(\mathbf{\hat{u}})\quad\mathrm{s.t.}\,\begin{cases}\|\mathbf{\Phi}\mathbf{\hat{u}}-\mathbf{v}\|_{2}\leq\varepsilon,\\ \mathbf{\hat{u}}\in\left[0,\,1\right]^{K},\end{cases}$ (32)
@@
[…]
@@
4:  $\overline{\mathbf{w}}_{n}^{(1)}\leftarrow\mathbf{w}_{n}^{(1)}+\gamma_{2}\mathbf{\Phi}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
5:  $\overline{\mathbf{w}}_{n}^{(2)}\leftarrow\mathbf{w}_{n}^{(2)}+\gamma_{2}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
6:  $\mathbf{w}_{n+1}^{(1)}\leftarrow\overline{\mathbf{w}}_{n}^{(1)}-\gamma_{2}P_{B^{\mathbf{v}}_{2,\,\varepsilon}}(\frac{1}{\gamma_{2}}\overline{\mathbf{w}}_{n}^{(1)})$;
7:  $\mathbf{w}_{n+1}^{(2)}\leftarrow\overline{\mathbf{w}}_{n}^{(2)}-\gamma_{2}P_{[0,\,1]^{K}}(\frac{1}{\gamma_{2}}\overline{\mathbf{w}}_{n}^{(2)})$;
8:end while
@@ III-B2 Image Restoration under Poisson Noise
In this observation model, the data-fidelity is expressed as the GKL divergence (refer to [65]), and image restoration can be reduced to the following optimization problem:
$\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\lambda\mathrm{GKL}_{\mathbf{v}}(\mathbf{\Phi}\mathbf{\hat{u}})+\mathcal{R}_{J}(\mathbf{\hat{u}})\quad\mathrm{s.t.}\,\mathbf{\hat{u}}\in\left[0,\,1\right]^{K}$ (35)
@@
[…]
@@
4:  $\overline{\mathbf{w}}_{n}^{(1)}\leftarrow\mathbf{w}_{n}^{(1)}+\gamma_{2}\mathbf{\Phi}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
5:  $\overline{\mathbf{w}}_{n}^{(2)}\leftarrow\mathbf{w}_{n}^{(2)}+\gamma_{2}(2\mathbf{u}_{n+1}-\mathbf{u}_{n})$;
6:  $\mathbf{w}_{n+1}^{(1)}\leftarrow\overline{\mathbf{w}}_{n}^{(1)}-\gamma_{2}\mathrm{prox}_{\frac{\lambda}{{\gamma_{2}}}\mathrm{GKL}}(\frac{1}{\gamma_{2}}\overline{\mathbf{w}}_{n}^{(1)})$;
7:  $\mathbf{w}_{n+1}^{(2)}\leftarrow\overline{\mathbf{w}}_{n}^{(2)}-\gamma_{2}{P_{[0,\,1]^{K}}}(\frac{1}{\gamma_{2}}\overline{\mathbf{w}}_{n}^{(2)})$;
8:end while
@@ III-C1 Convergent PnP-FBS [18]
[…]
@@ III-C2 Convergent PnP-ADMM [25, 24]
Several works have proposed PnP algorithms based on ADMM (PnP-ADMM) and established their convergence guarantees. Nair et al. have proposed a convergent PnP-ADMM and provided extensive analysis on its fixed-point [24]. Sun et al. have proposed a convergent PnP-ADMM algorithm that handles multiple data-fidelity terms, making it practical for large scale datasets [25]. In both studies, the convergence properties are ensured under the firm nonexpansiveness of the denoiser. PnP-ADMM can address a wider range of tasks than PnP-FBS, including image restoration with Poisson noise.
@@ III-C3 Convergent PnP-PDS[27]
We show the relationship between this algorithm and our PnP-PDS algorithm in Fig. 1. In essence, the algorithm in [27] is a special case of our Algorithm 1 without the box constraint.
@@ III-C4 Regularization by Denoising (RED)[45, 52]
$),\displaystyle\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\lambda\|\mathbf{\Phi}\mathbf{\hat{u}}-\mathbf{v}\|^{2}_{2}+\mathbf{\hat{u}}^{\top}(\mathbf{\hat{u}}-\mathfrak{G}(\mathbf{\hat{u}}$)), (43)
$),\displaystyle\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\lambda\mathrm{GKL}_{\mathbf{v}}(\hat{\mathbf{u}})+\mathbf{\hat{u}}^{\top}(\mathbf{\hat{u}}-\mathfrak{G}(\mathbf{\hat{u}}$)), (44)
@@
$).\displaystyle\nabla(\mathbf{\hat{u}}^{\top}(\mathbf{\hat{u}}-\mathfrak{G}(\mathbf{\hat{u}})))=2(\hat{\mathbf{u}}-\mathfrak{G}(\mathbf{\hat{u}}$)). (45)
@@
However, RED also has a considerable difficulty: the assumptions for the denoiser. First, local homogeneity is an unrealistic property, especially for deep denoisers. In fact, the gradients of the regularization terms computed in the RED algorithms often contain substantial computational errors due to the lack of local homogeneity in the denoisers [52]. Second, cyclically firm nonexpansiveness [46] is even stronger assumption than the firm nonexpansiveness assumed in our work. These challenges make it difficult to ensure the convergence of RED algorithms when we employ state-of-the-art denoisers, including DnCNN.
@@ IV Experiments
Method Algorithm Regularization
PnP-FBS FBS FNE* DnCNN
PnP-ADMM ADMM FNE* DnCNN
PnP-PDS (Unstable) PDS DnCNN (not FNE*)
PnP-PDS (NoBox) PDS FNE* DnCNN
TV PDS TV
RED (Gaussian) SD FNE* DnCNN
RED (Poisson) PDS FNE* DnCNN
DRUNet HQS DRUNet
BPG BPG Bregman Score Denoiser
Proposed PDS FNE* DnCNN
*FNE: firmly nonexpansive    SD: steepest descent
HQS: half-quadratic splitting
@@ IV-A Experimental Setup
$0.2246$
$0.1933$
$0.1907$
$0.1778$
$0.2255$
$0.2163$
$0.1917$
$0.1737$
$0.1763$
$0.1429$
Figure 2: Blur kernels used in our experiments. (a)–(h) correspond to kernels (1)–(8) in[66], while (i) and (j) correspond to the “GaussianA” and “Square” setup in [67], respectively. The Frobenius norm of each kernel is shown in the second row.
In the two experiments, we considered two types of observation operators: blur and random sampling. For the blur case, we used circular convolution with the ten kernels (a)–(j) from [66, 67], visualized in Fig. 2, and we scaled them so that $\|\mathbf{\Phi}\|_{\mathrm{op}}=1$. In Sec. IV-B-1 we evaluated all ten kernels (a)–(j), whereas in the other sections we used kernel (a) as a representative case. For the random sampling case, $20\%$ of the image pixels were randomly selected and masked. The operator norm also satisfies $\|\mathbf{\Phi}\|_{\mathrm{op}}=1$ in this case.
[…]
@@
Noise level Method (a) (b) (c) (d) (e) (f) (g) (h) (i) (j) Average
$\sigma=0.0025$ PnP-FBS $31.28^{*}$ $30.42^{*}$ $30.28^{*}$ $30.11^{*}$ $31.76^{*}$ -- $30.14^{*}$ $30.02^{*}$ $28.15^{*}$ $27.75^{*}$ --
Proposed $\mathbf{37.27}$ $\mathbf{36.63}$ $\mathbf{35.68}$ $\mathbf{36.64}$ $\mathbf{36.69}$ $\mathbf{37.49}$ $\mathbf{35.43}$ $\mathbf{35.58}$ $\mathbf{29.70}$ $\mathbf{30.85}$ $\mathbf{35.20}$
$\sigma=0.005$ PnP-FBS $31.16^{*}$ $30.35^{*}$ $30.22^{*}$ $30.04^{*}$ $31.67^{*}$ -- $30.07^{*}$ $29.95^{*}$ $28.12^{*}$ $27.71^{*}$ --
Proposed $\mathbf{34.04}$ $\mathbf{33.46}$ $\mathbf{32.95}$ $\mathbf{33.37}$ $\mathbf{34.01}$ $\mathbf{34.46}$ $\mathbf{32.71}$ $\mathbf{32.70}$ $\mathbf{28.87}$ $\mathbf{29.33}$ $\mathbf{32.59}$
$\sigma=0.01$ PnP-FBS $30.58$ $29.98^{*}$ $29.95^{*}$ $29.69^{*}$ $31.24$ $31.26$ $29.76^{*}$ $29.60^{*}$ $28.00^{*}$ $27.56^{*}$ $29.76$
Proposed $\mathbf{30.97}$ $\mathbf{30.52}$ $\mathbf{30.50}$ $\mathbf{30.37}$ $\mathbf{31.55}$ $\mathbf{31.74}$ $\mathbf{30.20}$ $\mathbf{30.09}$ $\mathbf{28.16}$ $\mathbf{27.85}$ $\mathbf{30.20}$
$\sigma=0.02$ PnP-FBS $\mathbf{28.26}$ $\mathbf{27.92}$ $\mathbf{28.35}$ $\mathbf{27.66}$ $29.34$ $29.27$ $\mathbf{28.18}$ $\mathbf{27.94}$ $\mathbf{27.43}$ $\mathbf{26.80}$ $\mathbf{28.12}$
Proposed $28.15$ $27.70$ $28.14$ $27.41$ $\mathbf{29.36}$ $\mathbf{29.28}$ $28.15$ $27.73$ $27.07$ $26.39$ $27.94$
$\sigma=0.04$ PnP-FBS $25.55$ $24.96$ $25.99$ $24.34$ $26.98$ $26.43$ $25.79$ $25.09$ $\mathbf{26.41}$ $25.18$ $25.67$
Proposed $\mathbf{26.00}$ $\mathbf{25.73}$ $\mathbf{26.29}$ $\mathbf{25.58}$ $\mathbf{27.05}$ $\mathbf{26.80}$ $\mathbf{26.13}$ $\mathbf{25.94}$ $26.18$ $\mathbf{25.42}$ $\mathbf{26.11}$
As the algorithm did not converge for $\lambda$ values in the range $0.8\lambda_{\mathrm{opt}}$ to $1.25\lambda_{\mathrm{opt}}$, we present the results for $\lambda=1.99$.
TABLE III: PSNR [dB] Values for 10 Blur Kernels Across Different Noise Levels. The Symbol “--” Denotes That the Algorithm Diverged and PSNR Values Could not be Obtained. The Better Results are Highlighted in Bold.
In each experiment, we selected the appropriate methods from PnP-FBS [18], PnP-ADMM [25], PnP-PDS (Unstable) [26], PnP-PDS (NoBox) [27], TV [62], RED [45], DRUNet [38], and Bregman Proximal Gradient (BPG) [51], and compared them with the proposed method. The optimization algorithms and regularization for each method are summarized in Table II. Details of PnP-FBS, PnP-ADMM, and RED are given in Section III-C. PnP-PDS (NoBox) corresponds to the proposed method without the box constraint, with all other settings aligned. We adopted the stopping criterion proposed by the authors for BPG. For the other methods, a fixed iteration count was used. Specifically, in the Poisson noise case with $\eta\in\{1,\,2,\,10\}$, the number of iterations was fixed to $4800$ for deblurring and $12000$ for inpainting. In the other cases, the number of iterations was fixed at $1200$ for the blur case and $3000$ for the random sampling case. For DRUNet, we employed the half-quadratic splitting algorithm, and used the same fixed iteration counts as the other methods. In addition, we also evaluate DRUNet (8 iter), where DRUNet is executed with 8 iterations of the HQS algorithm, to examine its performance in the low-iteration regime commonly used in practice.
For all settings, we used the DnCNN [18] trained with the noise level $\sigma_{J}$, for the regularization in PnP-FBS, PnP-ADMM, PnP-PDS (NoBox), RED, and the proposed method. This DnCNN is learned as described in Section II-D to promote its firm nonexpansiveness. For a fair comparison with PnP-FBS, whose the optimal parameter depends on $\sigma_{J}$, we adopted the model distributed by Pesquet et al., trained with $\sigma_{J}=0.01$ in Section IV-B, and used the same setting in Section IV-C. In Section IV-D, we investigate the influence of $\sigma_{J}$ by examining two cases: (i) $\sigma_{J}$ fixed to one of $\{0.0075,\,0.01,\,0.05,\,0.1\}$ and (ii) $\sigma_{J}$ randomly sampled from a uniform distribution over [0, 0.1][0,\,0.1] for each training sample.
@@ IV-B1 Versatility of a Once-Trained Denoiser
We investigate the versatility of the once-trained denoiser with the fixed value of $\sigma_{J}$ in different algorithms. Since the theoretical convergence guarantees in this work are built upon PnP-FBS, let us confirm that the versatility is enhanced by extending PnP-FBS to PnP-PDS.
As discussed in Section III-C1, the proposed method handles $\ell_{2}$ data-fidelity constraint, whereas PnP-FBS handles $\ell_{2}$ data-fidelity term. Following (39), we set $\varepsilon$ in (32) as $\varepsilon=\alpha\varepsilon_{\mathrm{opt}}$ where $\varepsilon_{\mathrm{opt}}:=\sigma\sqrt{K}$ and evaluated the performance over $\alpha\in[0.8,\,1.25]$. For PnP-FBS, we followed the heuristic in (40) and set $\lambda=\zeta\lambda_{\mathrm{opt}}$ where $\lambda_{\mathrm{opt}}:=\sigma_{J}/(2\sigma\|h\|)$ and $\|h\|$ is the Frobenius norm of the blur kernel $h$. We evaluated the performance over $\zeta\in[0.8,\,1.25]$. When the algorithm diverged within this range, we instead used $\lambda=1.99$.
Table III shows the variation in restoration performance across different noise levels and blur kernels. At low noise levels and for kernels with small Frobenius norms, PnP-FBS diverges with $\lambda$ close to $\lambda_{\mathrm{opt}}$ as discussed in Section III-C1. For example, in the case of $\sigma=0.01$ and kernel (b), we obtain $\lambda_{\mathrm{opt}}\approx 2.587$ because $\sigma_{J}=0.01$ and $2\sigma\|h\|\approx 0.0039$. However, the convergence condition (41) requires $\lambda<2$, forcing the use of suboptimal parameter choices. This discrepancy stems from the dependence of $\lambda_{\mathrm{opt}}$ on $\sigma_{J}$.
In contrast, the proposed method exhibits reasonable performance across all noise levels. This consistency demonstrates the high versatility of the given denoiser, thanks to the decoupling of $\varepsilon_{\mathrm{opt}}$ and $\sigma_{J}$ (see also the discussion in Section III-C1). Moreover, we note that $\varepsilon_{\mathrm{opt}}$ is independent of the blur kernels, which enables stable performance across all kernels under a consistent parameter choice.
@@ IV-B2 Restoration Performance
Deblurring $0.82$ $0.86$ $0.92$ $0.96$ $1.00$
Inpainting $0.90$ $0.82$ $0.82$ $1.00$ $1.00$
TABLE IV: The Values of $\alpha$ Used in Each Setting for the Proposed Method.
@@
PnP-PDS (NoBox) [27] $\underline{37.23}$ $34.03$ $31.00$ $28.24$ $\mathbf{26.00}$ -- -- -- -- --
@@
RED [45] $34.11$ $33.38$ $30.89$ $\mathbf{28.30}$ $24.87$ -- -- -- -- --
DRUNet [38] $32.60$ $31.80$ $28.54$ -- -- $33.21$ $32.68$ $31.36$ $26.71$ --
DRUNet (8 iter) [38] $35.33$ $\mathbf{34.50}$ $\mathbf{31.74}$ $21.65$ $11.38$ $13.84$ $13.84$ $13.82$ $13.77$ $13.57$
PnP-PDS (Proposed) $\mathbf{37.24}$ $\underline{34.04}$ $\underline{31.01}$ $\underline{28.26}$ $\mathbf{26.00}$ $\mathbf{35.61}$ $\mathbf{35.25}$ $\mathbf{34.43}$ $\mathbf{33.04}$ $\mathbf{31.00}$
TABLE V: The Average PSNR [dB] Values for Deblurring and Inpainting on the 7 Images From ImageNet. The Best Value is Highlighted in Bold and the Second Best Value is Underlined.
Figure 3: The evolution of $\|\nabla Q(\mathbf{u}_{n})\|_{\mathrm{sp}}^{2}$ of PnP-PDS (NoBox), PnP-FBS, RED, and the proposed method for inpainting at $\sigma=0.01$. The values in each plot are computed as the mean over 7 images.
$\mathbf{u}_{1}$ $\mathbf{u}_{753}$ $\mathbf{u}_{755}$ $\mathbf{u}_{757}$ $\mathbf{u}_{759}$ $\mathbf{u}_{761}$ $\mathbf{u}_{1001}$ $\mathbf{u}_{3000}$
PnP-PDS
$3.285$
$2492.4$
Proposed
$3.327$
$1.043$
$1.063$
$1.039$
$1.068$
$1.036$
$1.036$
$1.035$
Figure 4: Visualization of $\mathbf{u}_{n}$ of PnP-PDS (NoBox) and the proposed method, along with the corresponding values of $\|\nabla Q(\mathbf{u}_{n})\|_{\mathrm{sp}}^{2}$ for each $\mathbf{u}_{n}$. For visualization, each pixel value is clipped to the range $[0,1]$. The symbol “--” indicates that the value of $\|\nabla Q(\mathbf{u}_{n})\|_{\mathrm{sp}}^{2}$ could not be computed. Since $\mathbf{u}_{1001}$ and $\mathbf{u}_{3000}$ of PnP-PDS (NoBox) contained NaN values in Python, they could not be visualized.
Figure 5: The evolution of PSNR [dB] and $c_{n}$ of PnP-PDS (Unstable), DRUNet, and the proposed method over iterations for deblurring at $\sigma=0.01$. Each plot in the graph corresponds to each image extracted from ImageNet [68].
@@
We compare the performance of the proposed method with other methods, namely PnP-FBS[18], PnP-PDS (Unstable)[26], PnP-PDS (NoBox) [27], TV[62], RED[45], and DRUNet [38]. For constrained formulations such as PnP-PDS (Unstable), PnP-PDS (NoBox), TV, and the proposed method, we set $\varepsilon=\alpha\varepsilon_{\mathrm{opt}}$ in the same way as in the previous section. In particular, we provide the values of $\alpha$ for the proposed method in Table IV. For existing methods, we adopted the recommended values if parameter settings were suggested by the authors. Otherwise, we performed a linear search over the balancing parameter.
@@
Table V summarizes the average PSNR values over seven images, obtained with the optimal parameters for each method. The proposed method demonstrates superior performance compared to PnP-FBS, even though both methods employ the same denoiser for regularization. This is attributed to the difference in the versatility of the denoiser, as mentioned in Section IV-B1. Moreover, we observe that DRUNet (8 iter) achieves higher PSNR values than the proposed method in deblurring at $\sigma=0.005$ and $\sigma=0.01$. However, the proposed method outperforms DRUNet for all cases in larger iterations because of the difference in stability of the algorithms, which we further examine later with their convergence curves. This instability can be attributed to the lack of a theoretical convergence guarantee, and this observation is consistent with the literature, for instance in [69].
In Table V, we also note that PnP-FBS, PnP-PDS (NoBox), and RED fail to converge, and thus PSNR values cannot be obtained for the inpainting. To clarify this point, recall that the denoiser was trained with the penalty in (17), which was designed to promote nonexpansiveness for low-noise images sampled according to (18). During the PnP iterations, however, the inputs to the denoiser can become highly noisy and may have characteristics that differ from those of the training images. In such cases, the denoiser may no longer maintain its firm nonexpansiveness, potentially amplifying instability through successive iterations.
For numerical verification, we present the evolution of $\|\nabla Q(\mathbf{u}_{n})\|_{\mathrm{sp}}^{2}$ at each iteration in Fig. 3. Although PnP-PDS (NoBox), PnP-FBS, RED, and the proposed method employ the same denoiser, the values of $\|\nabla Q(\mathbf{u}_{n})\|_{\mathrm{sp}}^{2}$ in PnP-PDS (NoBox), PnP-FBS, and RED exhibit unstable behavior and tend to diverge. In contrast, the proposed method consistently maintains values of $\|\nabla Q(\mathbf{u}_{n})\|_{\mathrm{sp}}^{2}$ that satisfy the condition given in (16). In Fig. 4, we further compare the visualization of $\mathbf{u}_{n}$ obtained by PnP-PDS (NoBox) in a diverging case with that of the proposed method. For PnP-PDS (NoBox), erratic regions appear in the image, and the surrounding areas are affected in the proceeding iterations, leading to divergence of $\mathbf{u}_{n}$. Although similar irregular regions are observed at $\mathbf{u}_{753}$ in the proposed method, $\mathbf{u}_{n}$ remains stable and those regions are effectively removed in the subsequent iterations.